OpenStudy (anonymous):

Determine whether the given event is mutually exclusive or inclusive. Then find the probability. A card is drawn from a standard deck of cards. P(queen or jack)

OpenStudy (kirbykirby):

\(P(Queen)=4/52=P(Jack)\) By the addition rule: \(P(Queen-or-Jack)=4/52+4/52=8/52=2/13\)

OpenStudy (lopezking1):

52 cards in a deck. 4 queens and 4 jacks. Total = 8 Probability = 8/52 15 balls in total. 10 white or black Probability = 10/15

OpenStudy (kirbykirby):

mutually exclusive means: \(P(A-or-B)=P(A)+P(B)\) because in general, \(P(A-or-B)=P(A)+P(B)-P(A ∩ B)\)

OpenStudy (kirbykirby):

we can't have 1 card being both queen and jack at the same time, so the two events can't occur together, and are therefore mutually exclusive